Circular No. 6188

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 O1

The following precise positions have been reported:

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 24.45582 18^{h}43^{m}12.^{s}72 −32^{o}08′39′′.5  Garradd
24.46032 18 43 12.50 −32 08 39.5  _{”}
24.46243 18 43 12.43 −32 08 39.9  _{”}
24.54746 18 43 08.80 −32 08 32.6  Yoshida
24.54914 18 43 08.72 −32 08 33.1  Kojima
24.55007 18 43 08.67 −32 08 32.9  Kobayashi
24.55363 18 43 08.51 −32 08 32.0  Kojima
24.55410 18 43 08.46 −32 08 32.2 11  Kushida
24.55858 18 43 08.34 −32 08 30.4  Kojima
24.56441 18 43 08.02 −32 08 30.3  Kushida
24.56688 18 43 07.92 −32 08 31.9  Kobayashi
24.57585 18 43 07.58 −32 08 30.6  Yoshida
24.58288 18 43 07.23 −32 08 30.1  Kobayashi
24.58478 18 43 07.14 −32 08 31.0 12.0 Yoshida

G. J. Garradd (Loomberah, N.S.W.). 0.25-m f/4.1 reflector + CCD. 5-sec

exposures show coma diameter 1′ and slight elongation to the north. S. Yoshida (Wakayama). 0.25-m f/6.3 Schmidt-Cassegrain reflector + CCD.

Communicated by S. Nakano, Sumoto, Japan. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6.0 reflector + CCD.

Comet has some condensation and seems extended toward the north; m_{1} = 11.5–12.0. Communicated by Nakano.

T. Kobayashi (Oizumi). 0.25-m f/4.4 reflector + CCD. Communicated by

Nakano. R. Kushida and Y. Kushida (Yatsugatake South Base Observatory). 0.25-

m f/6.3 Schmidt-Cassegrain reflector + CCD. Tail 43′′ long in p.a. 5^{o}. Communicated by Nakano.

Total-visual-magnitude and coma-diameter estimates: July 24.164 UT, m_{1} = 11.0, 1′ (A. Hale, Cloudcroft, NM, 0.41-m reflector); 24.28, 10.6, 1′.7 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector; possible tail visible to the north).

1995 July 24 Daniel W. E. Green